From past experience, the supervisor of a coffee shop has found that 80% of customers entering the shop will order coffee and 30% of customers entering the shop will order snacks. It was also found that 20% of customers entering the shop would order coffee and snacks together.

(a) What is the probability that a customer entering the shop will either order coffee or snacks?

(b) What is the probability that a customer entering the shop will not order coffee and he/she will not order snacks?

(c) What is the probability that a customer entering the shop will order snacks given that he/she has ordered coffee?

(d) Are ‘order coffee’ and ‘order snacks’ statistically independent? Explain briefly.
Answer #1

a) P(Either coffee or snacks) = P(Coffee) + P(Snacks) - P(Both) = 0.80 + 0.30 - 0.20 = 0.90

b) P(Not order coffee and not order snacks) = 1 - P(Either coffee or snacks) = 1 - 0.90 = 0.10

c) P(Order snacks | Ordered coffee) = P(Both) / P(Ordered coffee) = 0.20/0.80 = 0.25

d) Since P(Order snacks | Ordered coffee) P(Order snacks), the events are not independent.